OpenStudy (anonymous):

Find all solutions of the equation. 5 sin x - 8sqrt2 = 3 sin x- 7sqrt 2 A. x = (5pi)/4 + nπ or x = (7pi)/4 + nπ B. x = (pi)/4 + nπ or x = (3pi)/4 + nπ C. x = (pi)/4+ 2nπ or x = (3pi)/4 + 2nπ D. x = (5pi)/4 + 2nπ or x = (7pi)/4 + 2nπ I have figured out sin x=(sqrt2)/2 but how do I solve for x now?

OpenStudy (anonymous):

5 sin(x)-8 squrt 2=3 sin(x)-7 sqrt (2)

OpenStudy (anonymous):

evaluate like term

OpenStudy (anonymous):

\[5\sin x- 8\sqrt{2}= 3\sin x-7\sqrt{2}\]

OpenStudy (anonymous):

I know what sin x equals \[\sqrt{2} \div2\]

OpenStudy (anonymous):

5 sin(x)-3 sin(x)=8 sqrt(2)-7sqrt(2) see clearer?

OpenStudy (jiteshmeghwal9):

5sinx-3sinx=7sqrt2+8sqrt2 2sinx=15sqrt2

OpenStudy (jiteshmeghwal9):

sinx=15sqrt2/2

OpenStudy (jiteshmeghwal9):

can u solve it from here @sarah43 ???

OpenStudy (anonymous):

@jiteshmeghwal9 its not gonna be 7 +8. its -7+8 making it 1 so its \[\sqrt{2}\]

OpenStudy (anonymous):

I already solved it ppl. I dont know how to solve for the x I have